# Differentially Private Covariate Balancing Causal Inference

### v1

@@ A.7 Proof of Lemma 3
Under Assumptions 16, Hirano et al. [2003], Zhao [2019] demonstrates that
@@
By [Li et al., 2018, Theorem 2], the first term $\xi_{1}$ converges to zero. For $\xi_{2}$, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon)$ from the proof of Theorem 3.2. As the propensity score $e(x)$ are continuous and has bounded derivative, $\hat{e}(x)-e_{\tilde{\theta}^{(1)}}=O_{p}(1/np\epsilon)$ as well by the first order taylor expansion. For the third term $\xi_{3}$, $\tilde{V}_{\alpha,\beta}=\hat{V}_{\alpha,\beta}+\nu_{V},$ where $\nu_{V}\sim\mathrm{Lap}(1,\Delta_{V,\alpha,\beta}/\epsilon)$ and $\Delta_{V,\alpha,\beta}=\frac{1}{2n\eta C_{\eta,\alpha,\beta}}$ by Lemma 4. The Laplace noise is asymptotically negligible and decays much faster than the rate of $n^{-1/2}$, thus $\xi_{3}$ converges to zero as well, which proves the consistency of $\tilde{V}_{\alpha,\beta}^{(2)}$. Finally, the clamping post-processing (15) only serves to obtain the valid variance estimator by projecting the negative estimator to the positive upper bound of the estimator. Since the convergence rate of $\hat{V}_{\alpha,\beta}$, $O_{p}(n^{-1/2})$, is slower than the rate of the Laplace noise $\nu_{V}$, $O_{p}(n^{-1})$, the post-processing will not be applied when $n$ is large enough as the Laplace noise $\nu_{V}$ tends to zero. ∎
@@ Appendix B Sieve Logistic Regression
We adopt the logistic regression with a polynomial series of the covariates for the propensity score estimation. This model can be interpreted as the sieve estimator [Geman and Hwang, 1982]. The sieve estimators are a class of non-parametric estimators that use progressively more complex models to estimate an unknown function as more data becomes available.
@@
The sieve approach is particularly appealing when modeling the propensity score. Hirano et al. [2003] studied the semiparametric efficiency of the IPW estimator when the dimension of the regressors $\phi(x)$ is allowed to increase as the sample size $n$ grows. Their renowned results claim that this sieve IPW estimator is semiparametrically efficient for estimating the WATE. Zhao [2019] showed that the semiparametric efficiency still holds if the Bernoulli likelihood, the loss function that Hirano et al. [2003] used to estimate the propensity score, is replaced by the Beta family of scoring rules $G_{\alpha,\beta}$, $-1\leq\alpha,\beta\leq 0$ or essentially any strongly concave scoring rule.
Specifically, using sieve logistic regression, where the series expansion of the log-odds function captures the complexities of treatment assignment mechanisms, as demonstrated by Hirano et al. [2003]. By employing the sieve logistic series estimator, the propensity score is approximated as:
@@ Appendix C Privacy-Aware Rejection Sampler for the KNG Mechanism
In this section, we state the sampling algorithm we use in Algorithm 1 for private propensity score and a required lemma from Awan and Rao [2024]. When the score function $S_{n,\alpha,\beta}$ for the causal estimand of interest is strongly concave, we adopt the privacy-aware rejection sampling technique of Awan and Rao [2024] to obtain a private parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from (10). This algorithm is designed to obtain the exact samples from the unnormalized density $\exp(S_{D}(x))$, not relying on approximate sampling techniques such as Markov Chain Monte Carlo. For this method, we assume that for the unnormalized target density $\pi_{D}$, we have a normalized density $U_{D}(x)$ as well as a constant and $c_{U,D}$ such that for all $x$:
@@
    if $Y\leq\frac{\tilde{\pi}(X)}{c_{U}U(X)}$ and anyAccepted == FALSE then
@@
Algorithm 3 Privacy-aware rejection sampling via squeeze functions [Awan and Rao, 2024, Algorithm 1]
The following lemma from Awan and Rao [2024, Lemma 33] establishes the upper bound for KNG.
@@
The convexity parameter $\lambda$ can be obtained from Zhao [2019]. Specifically, given $\mathrm{E}(\nabla_{\theta}g_{\theta}(X))=E\{\phi(X)\phi(X)^{\top}\}=I_{d}$ from the construction of the sieve regression for $g_{\theta}(X)=\theta^{\top}\phi(X)$, it suffices to consider the following derivatives for $S_{n,\alpha,\beta}$: $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),1)=(1-e)G^{\prime\prime}(e)(l^{-1})^{\prime}(g)=e^{\alpha}(1-e)^{\beta+1}$, $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),0)=-eG^{\prime\prime}(e)(l^{-1})^{\prime}(g)=-e^{\alpha+1}(1-e)^{\beta}$, $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),1)=\alpha e^{\alpha}(1-e)^{\beta+2}-(\beta+1)e^{\alpha+1}(1-e)^{\beta+1}$, and $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),0)=-(\alpha+1)e^{\alpha+1}(1-e)^{\beta+1}+\beta e^{\alpha+2}(1-e)^{\beta}.$ Therefore, for the estimation of ATE with $\alpha=\beta=-1$ for example, we have $\lambda=\frac{n\eta}{1-\eta}$ by Assumption 1.
@@
Awan and Rao [2024] also established a lower bound for KNG which can be used to ensure that the runtime does not depend on the database, at the cost of slower computation time. In our setting, we assume that the runtime is not available to the statistician.
@@ Appendix D Simulation details
Our methodology involves three key hyperparameters: $p\in(0,1)$, $10<q_{j}<$ 1 for $j=1,\ldots,4$ with $\sum_{j=1}^{4}q_{j}=1$, and $r\in(0,1)$. The parameters $p$ and $q_{j}$ govern the privacy level of the point estimator, specifically controlling the allocation of the privacy budget between the first and second stages of privatization. When there is no prior information to guide the choice of these parameters, it is generally recommended to assign equal weights to each privatization. Specifically, we set $p=0.2$ and $q_{1}=\ldots=q_{4}=0.25$ as the default choice.
According to Guha and Reiter [2024], we should choose the smallest feasible $M$ for their method, ensuring that the standard deviation of the Laplace noise from the subsampling and aggregation process remains significantly smaller than the sensitivity of the estimated variance of the WATE estimate obtained from the full data [Guha and Reiter, 2024], suggesting a fixed small value of $M$, e.g., $M=100$. However, we found that their method is sensitive to the choice of $M$ and works even better when $M$ grows with $n$. This observation aligns with their asymptotic results, where their estimator converges with $O_{p}(M^{-1})$. Therefore, we chose $M=\sqrt{N}$ instead of a fixed $M$, following their choice in their simulation.
Appendix E Covariate Balance Check for LaLonde [1986] Data under Privacy

### v2

@@ A.7 Proof of Lemma 3
Under Assumptions 16, Hirano et al. (2003); Zhao (2019) demonstrates that
@@
By (Li et al., 2018, Theorem 2), the first term $\xi_{1}$ converges to zero. For $\xi_{2}$, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon)$ from the proof of Theorem 3.2. As the propensity score $e(x)$ are continuous and has bounded derivative, $\hat{e}(x)-e_{\tilde{\theta}^{(1)}}=O_{p}(1/np\epsilon)$ as well by the first order taylor expansion. For the third term $\xi_{3}$, $\tilde{V}_{\alpha,\beta}=\hat{V}_{\alpha,\beta}+\nu_{V},$ where $\nu_{V}\sim\mathrm{Lap}(1,\Delta_{V,\alpha,\beta}/\epsilon)$ and $\Delta_{V,\alpha,\beta}=\frac{1}{2n\eta C_{\eta,\alpha,\beta}}$ by Lemma 4. The Laplace noise is asymptotically negligible and decays much faster than the rate of $n^{-1/2}$, thus $\xi_{3}$ converges to zero as well, which proves the consistency of $\tilde{V}_{\alpha,\beta}^{(2)}$. Finally, the clamping post-processing (15) only serves to obtain the valid variance estimator by projecting the negative estimator to the positive upper bound of the estimator. Since the convergence rate of $\hat{V}_{\alpha,\beta}$, $O_{p}(n^{-1/2})$, is slower than the rate of the Laplace noise $\nu_{V}$, $O_{p}(n^{-1})$, the post-processing will not be applied when $n$ is large enough as the Laplace noise $\nu_{V}$ tends to zero. ∎
@@ Appendix B Sieve Logistic Regression
We adopt the logistic regression with a polynomial series of the covariates for the propensity score estimation. This model can be interpreted as the sieve estimator (Geman and Hwang, 1982). The sieve estimators are a class of non-parametric estimators that use progressively more complex models to estimate an unknown function as more data becomes available.
@@
The sieve approach is particularly appealing when modeling the propensity score. Hirano et al. (2003) studied the semiparametric efficiency of the IPW estimator when the dimension of the regressors $\phi(x)$ is allowed to increase as the sample size $n$ grows. Their renowned results claim that this sieve IPW estimator is semiparametrically efficient for estimating the WATE. Zhao (2019) showed that the semiparametric efficiency still holds if the Bernoulli likelihood, the loss function that Hirano et al. (2003) used to estimate the propensity score, is replaced by the Beta family of scoring rules $G_{\alpha,\beta}$, $1\leq\alpha,\beta\leq 0$ or essentially any strongly concave scoring rule.
Specifically, using sieve logistic regression, where the series expansion of the log-odds function captures the complexities of treatment assignment mechanisms, as demonstrated by Hirano et al. (2003). By employing the sieve logistic series estimator, the propensity score is approximated as:
@@ Appendix C Privacy-Aware Rejection Sampler for the KNG Mechanism
In this section, we state the sampling algorithm we use in Algorithm 1 for private propensity score and a required lemma from Awan and Rao (2024). When the score function $S_{n,\alpha,\beta}$ for the causal estimand of interest is strongly concave, we adopt the privacy-aware rejection sampling technique of Awan and Rao (2024) to obtain a private parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from (10). This algorithm is designed to obtain the exact samples from the unnormalized density $\exp(S_{D}(x))$, not relying on approximate sampling techniques such as Markov Chain Monte Carlo. For this method, we assume that for the unnormalized target density $\pi_{D}$, we have a normalized density $U_{D}(x)$ as well as a constant and $c_{U,D}$ such that for all $x$:
@@
Algorithm 3 Privacy-aware rejection sampling via squeeze functions (Awan and Rao, 2024, Algorithm 1)
@@
    if $Y\leq\frac{\tilde{\pi}(X)}{c_{U}U(X)}$ and anyAccepted = FALSE then
@@
The following lemma from Awan and Rao (2024, Lemma 33) establishes the upper bound for KNG.
@@
The convexity parameter $\lambda$ can be obtained from Zhao (2019). Specifically, given $\mathrm{E}(\nabla_{\theta}g_{\theta}(X))=E\{\phi(X)\phi(X)^{\top}\}=I_{d}$ from the construction of the sieve regression for $g_{\theta}(X)=\theta^{\top}\phi(X)$, it suffices to consider the following derivatives for $S_{n,\alpha,\beta}$: $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),1)=(1-e)G^{\prime\prime}(e)(l^{-1})^{\prime}(g)=e^{\alpha}(1-e)^{\beta+1}$, $\frac{d}{dg}S_{n,\alpha,\beta}(l^{-1}(g),0)=-eG^{\prime\prime}(e)(l^{-1})^{\prime}(g)=-e^{\alpha+1}(1-e)^{\beta}$, $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),1)=\alpha e^{\alpha}(1-e)^{\beta+2}-(\beta+1)e^{\alpha+1}(1-e)^{\beta+1}$, and $\frac{d^{2}}{dg^{2}}S_{n,\alpha,\beta}(l^{-1}(g),0)=-(\alpha+1)e^{\alpha+1}(1-e)^{\beta+1}+\beta e^{\alpha+2}(1-e)^{\beta}.$ Therefore, for the estimation of ATE with $\alpha=\beta=-1$ for example, we have $\lambda=\frac{n\eta}{1-\eta}$ by Assumption 1.
@@
Awan and Rao (2024) also established a lower bound for KNG which can be used to ensure that the runtime does not depend on the database, at the cost of slower computation time. In our setting, we assume that the runtime is not available to the statistician.
@@ Appendix D Simulation details
According to Guha and Reiter (2024), we should choose the smallest feasible $M$ for their method, ensuring that the standard deviation of the Laplace noise from the subsampling and aggregation process remains significantly smaller than the sensitivity of the estimated variance of the WATE estimate obtained from the full data (Guha and Reiter, 2024), suggesting a fixed small value of $M$, e.g., $M=100$. However, we found that their method is sensitive to the choice of $M$ and works even better when $M$ grows with $n$. This observation aligns with their asymptotic results, where their estimator converges with $O_{p}(M^{-1})$. Therefore, we chose $M=\sqrt{N}$ instead of a fixed $M$, following their choice in their simulation.
Appendix E Addtional analyses
E.1 Covariate Balance Check for LaLonde (1986) Data under Privacy